# A Graph-Theoretical Perspective on Law Design for Multiagent Systems

### v1

@@ A.3 Proof of Lemma 3
($\Rightarrow$) Suppose $L$ is gap-free but not useful. Then, $\mathbb{P}^{L}\neq\varnothing$ by Lemma 1. Consider an arbitrary profile $\delta\in\mathbb{P}^{L}$. Then, $\delta\in\mathbb{P}$ and
@@
Consider an arbitrary agent $a\in\mathcal{A}$. Then, $\delta_{a}\notin L$ by item 1 of Definition 8 and statement (3). Thus, by statement (1), Definition 3, and the arbitrariness of $a\!\in\!\mathcal{A}$,
@@
and $\delta\in\mathbb{P}\cap\prod\Delta^{L}=\mathbb{P}^{L}$. The latter implies $\mathbb{P}^{L}\neq\varnothing$. Hence, law $L$ is not useful by Lemma 1. Meanwhile, by statements (3), (4), and item 2 of Definition 8, there is no safe action of (the arbitrary) agent $a$ in the law-imposed game $(\mathcal{A},\Delta^{L},\mathbb{P}^{L})$. ∎
@@ B.1 Proof of Theorem 2
To increase the readability of the proof of Theorem 2, let us first prove the next lemma.
@@ B.2 Proof of Theorem 3
As a standard notation in mathematics,
@@ C.4 Proof of Lemma 5
$(\Leftarrow)$ Suppose $d\in\Delta_{a}^{L}$ and $L$ is a vertex cover in the graph $\mathcal{H}^{a,d}_{(\mathcal{A},\Delta,\mathbb{P})}$. Then, $L\cap(\mathcal{S}(\delta)\setminus\{d\})\neq\varnothing$ for each profile $\delta\in\mathbb{P}$ such that $\delta_{a}=d$ by item 2 of Definition 18 and item 1 of Definition 13. Thus, $L\cap\mathcal{S}(\delta)\neq\varnothing$ for each profile $d\in\mathbb{P}$ such that $\delta_{a}=d$. Hence, $d$ is a safe action of agent $a$ in the law-imposed game $(\mathcal{A},\Delta^{L},\mathbb{P}^{L})$ by Lemma 2 and the $(\Leftarrow)$ part assumption $d\in\Delta_{a}^{L}$. ∎
@@ C.5 Proof of Theorem 9
If item 2 is false, then, by item 2 of Definition 13, there is an agent $a$ and a safable action $d\in\Delta_{a}^{L}$, and a vertex cover $L^{\prime}$ in the graph $\mathcal{H}^{a,d}_{(\mathcal{A},\Delta,\mathbb{P})}$ such that

### v2

@@ A.3 Proof of Lemma 3
($\Rightarrow$) Suppose $L$ is gap-free but not useful. Then, $\mathbb{P}^{L}\neq\varnothing$ by Definition 4. Consider an arbitrary profile $\delta\in\mathbb{P}^{L}$. Then, $\delta\in\mathbb{P}$ and
@@
Consider an arbitrary agent $a\in\mathcal{A}$. Then, $\delta_{a}\notin L$ by item 1 of Definition 8 and statement (3). Thus, by statement (1) and the arbitrariness of $a$,
@@
Hence, law $L$ is not useful by Lemma 1 and that $\delta\in\mathbb{P}$ in statement (3). Meanwhile, by statements (3), (4), and item 2 of Definition 8, there is no safe action of (the arbitrary) agent $a$ in the law-imposed game $(\mathcal{A},\Delta^{L},\mathbb{P}^{L})$. ∎
@@ B.1 Proof of Theorem 2
Let us first illustrate Definition 14 with an example as illustrated in Figure 1. In this $3$-graph, there are five vertices $r,s,t,u,v$ and three edges: $e_{1}=\{s\}$, $e_{2}=\{t,u\}$, and $e_{3}=\{r,u,v\}$. The induced game, by item 1 of Definition 14, is played within three agents: $1,2,3$. Each agent, by item 2 of Definition 14, has access to five actions: $r,s,t,u,v$. Moreover, by item 3 of Definition 14, there are three prohibited outcomes: $\delta^{e_{1}}$, $\delta^{e_{2}}$, $\delta^{e_{3}}$. In particular, taking the predefined order in item 3 of of Definition 14 as the lexicographical order. Then, $\delta^{e_{1}}_{1}=\delta^{e_{1}}_{2}=\delta^{e_{1}}_{3}=s$; $\delta^{e_{2}}_{1}=\delta^{e_{2}}_{3}=u,\delta^{e_{2}}_{2}=t$; $\delta^{e_{3}}_{1}=u,\delta^{e_{3}}_{2}=v,\delta^{e_{3}}_{3}=r$. The induced game is as illustrated in Figure 2.
Figure 1: A sample of 33-graph.
Figure 2: The game induced from the $3$-graph in Figure 1 by Definition 14.
Observe Figure 3 that the set $\{s,u\}$ is a vertex cover of the graph in Figure 1. If a law bans these two actions in the induced game in Figure 2, then, as illustrated in Figure 4, all outcomes corresponding to the greyed cells become unlawful, including all three prohibited outcomes. In other words, all lawful profiles are non-prohibited. Therefore, the set $\{s,u\}$ is also a useful law in this game.
Figure 3: Illustration of a vertex cover in the $3$-graph in Figure 1.
Figure 4: Illustration of a useful law in the game in Figure 2.
More generally, one can prove Theorem 2 below, which bridges a vertex cover and a useful law. To increase the readability of its proof, let us first prove the next lemma.
@@ B.2 Proof of Theorem 3
First, note that, as a standard notation in mathematics,
@@
We use the introductory example to illustrate an induced graph $(\bigcup\Delta,\mathcal{S}(\delta))$. Note that the game in each three-day cycle of this example can be captured in Figure 5. In this game, there are three agents $a$, $b$, and $c$; each agent can choose to dump on the first, the second, or the third day in a three-day cycle. The bio-hazard symbol ☣  captures the outcomes where the fish is killed, which happens when all three factories dump on the same day. This setting contains $3\times 3$ relavant actions: $d_{a}^{1}$, $d_{a}^{2}$, $d_{a}^{3}$, $d_{b}^{1}$, $d_{b}^{2}$, $d_{b}^{3}$, $d_{c}^{1}$, $d_{c}^{2}$, $d_{c}^{3}$.
Figure 5: Matrix game in accordance with the introductory example.
The induced graph consists of nine vertices: $d_{a}^{1}$, $d_{a}^{2}$, $d_{a}^{3}$, $d_{b}^{1}$, $d_{b}^{2}$, $d_{b}^{3}$, $d_{c}^{1}$, $d_{c}^{2}$, $d_{c}^{3}$, as well as three edges: $e_{1}=\{d_{a}^{1},d_{b}^{1},d_{c}^{1}\}$, $e_{2}=\{d_{a}^{2},d_{b}^{2},d_{c}^{2}\}$, and $e_{3}=\{d_{a}^{3},d_{b}^{3},d_{c}^{3}\}$. The induced graph is illustrated in Figure 6.
Figure 6: The graph induced from the game in Figure 5.
Now, we formally prove that the tuple $(\bigcup\Delta,\mathcal{S}(\mathbb{P}))$ is an $|\mathcal{A}|$-graph.
@@ B.4 Proof of Theorem 4
Observe Figure 7 that the law $L_{0}$ in the introductory example is a vertex cover in the graph in Figure 6.
Figure 7: $L_{0}$ is a useful law in the game in Figure 5 and a vertex cover in the graph in Figure 6.
More generally, we can prove the next theorem.
@@ B.5 Proof of Theorem 5
Figure 8: An illustration of Definition 15 that uses the graph in Figure 6.
Note that, by Theorem 4, reducing a useful law corresponds to finding a smaller vertex cover within an existing one. For instance, the law $L_{0}$ forms a vertex cover in the graph in Figure 6, as illustrated in Figure 7. To reduce law $L_{0}$ into a useful law is to select a subset of $L_{0}$ that can cover the graph in Figure 6. Observe that, this is equivalent to obtain a vertex cover in the graph shown in Figure 8. In other words, the smaller vertex cover can be regarded as a vertex cover in a subgraph induced by the original cover, which is formalised as the theorem below.
@@ C.1 Proof of Theorem 6
Definition 16 indeed characterises a polynomial algorithm to construct the game $(\bar{\mathcal{A}},\bar{\Delta},\bar{\mathbb{P}})$ for any given game $(\mathcal{A},\Delta,\mathbb{P})$. To illustrate this definition, let us use the game in Figure 5 as the input game $(\mathcal{A},\Delta,\mathbb{P})$, and construct the output game $(\bar{\mathcal{A}},\bar{\Delta},\bar{\mathbb{P}})$, as illustrated in Figure 9.
Figure 9: An illustration of Definition 16 using the game in Figure 5.
In this output game, an additional agent $\gamma$ is introduced alongside the original three agents $a,b,c$ (item 1 of Definition 16). Each of the original agents $a,b,c$ has one more available action $n$, in addition to their original actions, while the newly introduced agent $\gamma$ has two available actions $p$ and $n$ (item 2 of Definition 16). If agent $\gamma$ chooses action $p$ (see the above part of Figure 9), then outcomes are prohibited only when the behaviours of the original agents constitute a prohibited outcome in the input game, as labelled by ☣1 in Figure 9 (the first bullet in item 3 of Definition 16). If agent $\gamma$ chooses action $n$ (see the below part of Figure 9), then outcomes are prohibited when exactly one of the original agents chooses their original actions, as labelled by ☣2 in Figure 9 (the second bullet in item 3 of Definition 16), or when every original agent chooses action $n$, as labelled by ☣3 in Figure 9 (the third bullet in item 3 of Definition 16).
@@
Recall Lemma 1 that a useful law $L$ in the game $(\mathcal{A},\Delta,\mathbb{P})$ intersects with $\mathcal{S}(\delta)$ for each profile $\delta\in\mathbb{P}$. Then, by item 1 of the above lemma, the law $L$ intersects with $\mathcal{S}(\bar{\delta})$ for each profile $\bar{\delta}\in\bar{\mathbb{P}}_{1}$. Also observe that $p,n\notin L$ and agent $\gamma$ takes action $n$ in each profile in the set $\bar{\mathbb{P}}_{2}\cup\bar{\mathbb{P}}_{3}$. Then, by Lemma 2, the law $L$ makes $p$ a safe action of agent $\gamma$ in the game $(\bar{\mathcal{A}},\bar{\Delta}^{L},\bar{\mathbb{P}}^{L})$, which means that the law $L$ is gap-free in the game $(\bar{\mathcal{A}},\bar{\Delta},\bar{\mathbb{P}})$ by Lemma 3. Theorem 6 below formalises the above observation and establishes its converse as well.
@@ C.4 Proof of Lemma 5
To illustrate Definition 18, let us use the game in Figure 9. Consider the action $p$ of agent $\gamma$, where $p$ is safable by Lemma 4. The induced graph by Definition 18 is visualised in Figure 10. In particular, by item 1 of Definition 18, the graph consists of ten vertices corresponding to the ten (out of eleven, except for action $p$) actions in Figure 9. Meanwhile, there are three prohibited outcomes $\delta$ such that $\delta_{\gamma}=p$ in Figure 9, as labelled by the sign ☣1. Accordingly, by item 2 of Definition 18, there are three edges in Figure 10, each corresponding to a ☣1-labelled outcome in Figure 9. For instance, edge $e_{1}=\{d_{a}^{1},d_{b}^{1},d_{c}^{1}\}$ in Figure 10 corresponds to the left-most ☣1-labelled outcome in Figure 9, which consists of the actions $p,d_{a}^{1},d_{b}^{1}$, and $d_{c}^{1}$. The same is true for edge $e_{2}$ in Figure 10 and the middle ☣1-labelled outcome in Figure 9, as well as edge $e_{3}$ in Figure 10 and the right-most ☣1-labelled outcome in Figure 9.
Figure 10: An illustration of Definition 18 using the the game in Figure 9, agent $\gamma$, and action $p$.
Observe that Definition 18 indeed implies a polynomial algorithm to convert a game $(\mathcal{A},\Delta,\mathbb{P})$, an agent $a\in\mathcal{A}$, and an action $d\in\Delta_{a}$ into the graph $\mathcal{H}^{a,d}_{(\mathcal{A},\Delta,\mathbb{P})}$. The next lemma formally captures the aforementioned observation about their relation and thereby bridges “a safe action in a law-imposed game” with a vertex cover in the induced game.
@@
$(\Leftarrow)$ Suppose $d\in\Delta_{a}^{L}$ and $L$ is a vertex cover in the graph $\mathcal{H}^{a,d}_{(\mathcal{A},\Delta,\mathbb{P})}$. Then, $L\cap(\mathcal{S}(\delta)\setminus\{d\})\neq\varnothing$ for each profile $\delta\in\mathbb{P}$ such that $\delta_{a}=d$ by item 2 of Definition 18 and item 1 of Definition 13. Thus, $L\cap\mathcal{S}(\delta)\neq\varnothing$ for each profile $\delta\in\mathbb{P}$ such that $\delta_{a}=d$. Hence, $d$ is a safe action of agent $a$ in the law-imposed game $(\mathcal{A},\Delta^{L},\mathbb{P}^{L})$ by Lemma 2 and the $(\Leftarrow)$ part assumption $d\in\Delta_{a}^{L}$. ∎
By the above lemma, for a law $L$ in the game in Figure 9, action $p$ is a safe action of agent $\lambda$ under law $L$ if and only if $L$ is a vertex cover in the graph in Figure 10. This property, however, is easily observable through another chain of equivalence. Observe that a vertex cover in the graph in Figure 10 is exactly a vertex cover in the graph in Figure 6, the latter of which, by Theorem 4, is a useful law in the game in Figure 5. Such a useful law, by Theorem 6 and its proof, is gap-free in the game in Figure 9 by rendering action $p$ a safe action of agent $\gamma$ in the law-imposed game. This observation provides insight into the construction presented in Definition 16.
@@ C.5 Proof of Theorem 9
If item 2 is false, then, by item 2 of Definition 13, there is an agent $a$, a safable action $d\in\Delta_{a}^{L}$, and a vertex cover $L^{\prime}$ in the graph $\mathcal{H}^{a,d}_{(\mathcal{A},\Delta,\mathbb{P})}$ such that